I can not figure out this problem, please help: In most businesses, increasing prices of their product can have a negative effect on the number of customers of the business. A bus company in a small town has an average number of riders of 800 per day. The bus company charges $2.25 for a ride. They conducted a survey of their customers and found that they will lose approximately 40 customers per day for each $.25 increase in fare.
Answer by checkley71(8403)

THE FORMULA FOR THIS PROBLEM IS 225+25X=800-40X BUT WHAT ARE YOU SEEKING?
WHEN DOES THE COMPANY GET TO 0 CUSTOMERS? ANS. $7.25
HOW MANY CUSTOMERS DO THEY LOSE IF THEY INCREASE THE FARE 100? 800-160=540